Each of Basit’s friends added the number of the day and the number of the month of his birthday and
obtained 35. Their birthdays are all different. What is the maximum possible number of Basit’s friends?

Answer by ikleyn(52852)

.
December,  23    gives  (12 + 23 = 35)

November,  24    gives  (11 + 24 = 35)

October,   25    gives  (10 + 25 = 35)

September, 26    gives  ( 9 + 26 = 35)

August,    27    gives  ( 8 + 27 = 35)

July,      28    gives  ( 7 + 28 = 35)

June,      29    gives  ( 6 + 29 = 35)

May,       30    gives  ( 5 + 30 = 35)


and that's all.


ANSWER.   Maximum possible number of Basit’s friends is equal to the number of lines in my list  ( 8 ).

Solved.